A project has the following cash flows Year Cash Flow 0 $17,400 8,100 9,400 7,900 2 What is the NPV at a discount rate of zero percent? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.) NPV What is the NPV at a discount rate of 11 percent? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) NPV What is the NPV at a discount rate of 23 percent? (Negative amount should be indicated by a minus sign. Do not round intermediate calculations and round your answer to 2 decimal places, e.g. 32.16.) NPV What is the NPV at a discount rate of 29 percent? (Negative amount should be indicated by a minus sign. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) NPV
Answer #1

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

a.Present value of inflows=(8100+9400+7900)=$25400

NPV=Present value of inflows-Present value of outflows

=(25400-17400)=$8000.

b.Present value of inflows=8100/1.11+9400/1.11^2+7900/1.11^3

=$20702.96

NPV=$20702.96-$17400

=$3302.96

c.Present value of inflows=8100/1.23+9400/1.23^2+7900/1.23^3

=$17043.93

NPV=$17043.93-$17400

=$(356.07)(Negative)

d..Present value of inflows=8100/1.29+9400/1.29^2+7900/1.29^3

=$15607.85

NPV=$15607.85-$17400

=$(1792.15)(NEGATIVE)(APPROX).